Science Prep Night 20130526 UT

Summary

  • MODS working with heb_r power in bypass (see IT#4351)
  • LUCI scripting failing for instrument config problem that does not happen manually, Team notified
  • LBC seems fine, no problems.

LUCI

luci.20130526.NNNN.fits

Started with LUCI, but ran into MODS issue before opening.

Good seeing at start: 0.5-0.7"

Getting LUCI scripts failing for an instrument change. The failure (details below) seems to be associated with the grating, even though it initializes and moves OK when things are done manually. There is no error on the UI. I don't want to restart the SW because Peter has the MOS grabber warming up (plus, I did that as part of the checkout on Wednesday).

The script was generated with Dennis' script creator web page and it passes both lucichk4.26.pl and the LUCI script execution test mode. I've looked at it with a couple editors on both Mac and Linux and don't see any odd stray characters or anything like that.

From the LUCI message database: Details of Message: 7188715 MySQL Timestamp: 2013-05-26 04:02:36 Software Timestamp: 1369540956804 = 2013-05-26 04:02:37 Time Service Client: unsync Type: #3 = ERROR Level: #2 = MEDIUM Usage: #2 = Lucifer ONE Host: #2 = lucifer[192.168.0.1] Program: #28 = ObservationScript Address: Package: #7 = de.rub.astro.lucifer.operation.instrumentManager
Class
#77 = InstrumentManagerClient
Method
#69 = setNewInstrumenSetup
File
#77 = InstrumentManagerClient.java Line = 169
Content: [459] #3433 = Error setting new instrument setup from instrument manager server: rmi://localhost:1098/InstrumentManager Attachment: [44] #1124 = java.lang.Exception: RemoteException occurred in server thread; nested exception is: java.rmi.RemoteException: de.rub.astro.lucifer.control.TimeOutException: Command 2 timed out!

LUCI Team notified via email. I have heard back from Walter before the end of the night.

Configured instrument manually, commented out instrument section from script. Script ran OK.

sn2011fe_H (Milne) Files 0001-0040

Error setting filter wheel on new reconfigure. Init recovered.

noslitdither No mask in fpu, but does 20-points along a "slit" to check orientation

06:13 Switch to MODS...

Reconfigure LUCI for blind+blind filters, N3.75 camera, log out of UI.

Log out of LUCI UI.

LUCI:/data/luci partial clean up, now 46GB free.

Mask table verified, all .lms files exist at the specified location.

Still to do: Check that lamps work
Check darks (see: LBTtools/LUCI/readnoise script)
  • O2DCR 2.0s (usually around 11.6 e-, 12.6 for detector #2)
  • MER10 10.0s (usually around 5.5 e-, 7.75 for detector #2)

MODS

MODS1 IC restart for the red channel fails to complete. Red heb_r breaker fault (same as IT#4351 but other side). Tried shutting down and restarting SW but that did not work. Called Rick who confirmed it looked like IT#4351. It does not appear that we have a spare on the mountain. MODS is running in "bypass" mode. Email sent to telescopework to be careful if there are any glycol shutoffs since heb_r will overheat if not shut off.

Tests on Sommariva Field

BS9169 star @ 17hr to collimate on axis. Clear active optics and allow to collimate

n1625_r1319.img
n1625_r1572.img
n1625.img

Test on StoneM2 field
Clear active optics and run Collim_OnAxis.img (0008-11) then StoneM2reg_u60.img (0012-15)
Clear active optics and run Collim_OnAxis.img (0016-19) then StoneM2reg_u30.img (0020-23)
Clear active optics and run Collim_OnAxis.img (0024-27) then StoneM2reg.img (0028-31)
Clear active optics and run Collim_OnAxis.img (0032-33) then StoneM2reg_d30.img (0034-37)
Clear active optics and run Collim_OnAxis.img (0038-41) then StoneM2reg_d60.img (0042-45)
Clear active optics and run Collim_OnAxis.img (0046-47?) then StoneM2reg_d60.img (0048-51)

execMODS modsSleep.pro

/lhome/data had only 400MB free space left! Cleared 76GB.

LBC

Started reconfigure at start of twilight.

10:49 Started DOFPIA, /first, /X2 on BS9178 Both sides converged in 3 cycles. First cycle on both sides the pupil were small...adjust GOs in the collimation model (-4700 red -8800 blue)?

11:02 Preset to BS9177. DOFPIA, /X2 Collimated in 2 cycles, both sides.

Into twilight, closing up.

-- DavidThompson - 26 May 2013
Topic revision: r5 - 28 May 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback